Evaluate the expressions

2/5 divided by 3/10
-3m(5/6) divided by 1/6
-1/8 divided by 3(3/4)
3(1/5) divided by 5( -1/2 )

Answer by stanbon(48544)

2/5 divided by 3/10
2/5 *10/3 = 2/1 * 2/3 = 4/3
------------------
-3m(5/6) divided by 1/6
-3m(5/6) * 6/1 = -15m
---------------

-1/8 divided by 3(3/4)
-1/8 * 4/15 = -1/2 *1/15 = -1/30
----------------


3(1/5) divided by 5( -1/2 )
16/5 * -2/5= -32/25
Cheers,
Stan H.